OpenStudy (anonymous):

convert 598m^2 into ha

OpenStudy (anonymous):

round to two decimal places

OpenStudy (mertsj):

1 hectare = 10,000 m^2

OpenStudy (anonymous):

when i divide i get 16.722

OpenStudy (mertsj):

\[598 m^2 \times \frac{1 ha}{10000 m^2}=\frac{598ha}{10,000}=.0598 ha=.06 ha\]

OpenStudy (anonymous):

so im just rounding the first number which would be 5 into 6?

OpenStudy (mertsj):

doesn't it say round to the nearest hundredth?

OpenStudy (anonymous):

it does its just worded different lol